OpenStudy (anonymous):

The third angle in an isosceles triangle is 1/22 as large as each of the two base angles. Find the measure of each angle.

OpenStudy (wolf1728):

Base angle = x third angle = 1/22x x + x + 1/22x = 180 x + x + (1/22)x = 180 2x + (1/22)x= 180 2.0454545455 x = 180 x = 88 Both base angles = 88 and third angle equals (88/22) = 4 degrees. 88 + 88 + 4 = 180°